OpenStudy (anonymous):

What is the length of the hypotenuse, x, if (20, 21, x) is a Pythagorean triple?

OpenStudy (alexalas):

well a^2 + b^2=c^2 a=20 b=21 a Pythagorean triple all sides are whole numbers like 1,2,3,4,5,6,7,8,9,10....

OpenStudy (anonymous):

the choices are 22 29 41 42

OpenStudy (alexalas):

a^2 + b^2= ? find ?

OpenStudy (anonymous):

Answer and get a medal

OpenStudy (alexalas):

I am sorry I will not give direct answers. You give me the answer I tell you if you got it correct. Do not start guessing or I will leave. I want you to work it out so I know you know how to do it. okay?

OpenStudy (alexalas):

well a^2 + b^2=c^2 a=20 b=21 find c^2

OpenStudy (anonymous):

after you find a^2 And B^2 do you divide?

OpenStudy (anonymous):

i am confused.

OpenStudy (anonymous):

it would be 29?

OpenStudy (alexalas):

let me give you a few examples  3^2+4^2=25 the you square root square root is what times itself = this number so 5*5=25 so the square root of 25 is 5